OpenStudy (pottersheep):

How come the answer to this is "no solution"? x^2 +4x +1 = 0

OpenStudy (pottersheep):

I know it isnt factorable......but cant you plug it into the quadratic equation?

OpenStudy (precal):

This quadratic has 2 solutions: \[(-4\pm \sqrt{12})/2\] In general, quadratics have 3 possible solutions (depending on how many times the function crosses the x-axis) 1) one solution eg \[(x-3)^2=0\] the solution is x=3 (double root) 2) two solutions eg \[x^2-4=0\] the solutions are x=2 and x=-2 3) no solutions (look for the negative number underneath the square root symbol) and the function does not cross the x axis.

OpenStudy (pottersheep):

It has two solutions? ergh but my answer sheet ssays none! Thanks

OpenStudy (precal):

Look at the graph.

OpenStudy (pottersheep):

There is no graph :S I'll review my notes now

OpenStudy (precal):

Do you have a graphing calculator? You could use winplot (it is a free software that allows you to view graphs)

OpenStudy (pottersheep):

I dont have a graphing calculator but I'll go google up winplot ~~ Thanks again

OpenStudy (precal):

google online graphing calculator or something like that